Pls help in solving.
A bag contains 24 tennis balls,some white and some green.if a ball is chosen at random, the probability of getting a green ball is 3/8. How many white balls are there in the bag

Answer by jim_thompson5910(35256)

There are (3/8)*24 = 9 green balls. So there are 24-9 = 15 white balls

Alternative way to do it: 1-3/8 = 5/8 ---> (5/8)*24 = 15